Ultraviolet radiations of 6.2 eV falls on an aluminium surface. K.E. of fastest electron emitted is (work function = 4.2 eV)


        (a)    3.2 × 10-21 J


        (b)    3.2 × 10-19 J


        (c)    7 × 10-25 J


        (d)    9 × 10-32 J


Grade:11

1 Answers

Amit Saxena
10 years ago

(b)

K.E. of fastest electron

=      E – W0 = 6.2 - 4.2 = 2.0 eV

=      2 × 1.6 × 10-19 = 3.2 × 10-19 J